OpenStudy (anonymous):

For an object whose velocity in ft/sec is given by v(t) = −3t^2 + 5, what is its displacement, in feet, on the interval t = 0 to t = 2 secs? a: 6.607 b: 2 c: −2.303 d: 2.303

OpenStudy (anonymous):

@misty1212 can u plz help? :)

OpenStudy (anonymous):

I think its a, 6.607 but double-checking is good :)

OpenStudy (misty1212):

if i am not mistaken it is \[\int_0^2(-3t+5)dt\]

OpenStudy (solomonzelman):

no, this is correct. s(t) is integral of v(t)

OpenStudy (misty1212):

oh but i wrote it wrong doe

OpenStudy (solomonzelman):

yes, but conceptualized correctly:)

OpenStudy (misty1212):

\[\int_0^2(-3t^2 +5)dt\]

OpenStudy (misty1212):

your answer, however, is wrong

OpenStudy (anonymous):

kk integral...

OpenStudy (solomonzelman):

put it inline, \(\large\color{slate}{\displaystyle\int\limits_{0}^{2} \left( -3t^2+5\right)~~dt}\)

OpenStudy (solomonzelman):

`\(\large\color{slate}{\displaystyle\int\limits_{0}^{2} \left( -3t^2+5\right)~~dt}\)` (don't mean to give a latex class)

OpenStudy (anonymous):

? i can't see it

OpenStudy (misty1212):

that does make the numbers look better for sure

OpenStudy (misty1212):

\[\int\limits_{0}^{2}\left(-3t^2+5\right)dt\]

OpenStudy (misty1212):

thnx!

OpenStudy (anonymous):

thank u :) i can see that but do u mind walking me thru the integration process

OpenStudy (anonymous):

well i think its c, -2.303 @misty1212 :)